1) The angles of a pentagon are, x°, (x+20)°, (x-15)°, 2x°, (3x/2+30)°. Find x.


1
Expert's answer
2021-09-16T00:49:28-0400

Task #237198


n=5 (5 angles) 


x°+(x+20)°+(x-15)°+2x°+3x/2 +30°=180°•(n-2) 


5x°+3x/2=540°-20°+15°-30°


13x°=1010


x°=77,69